3.) Determine the following:

Ventricular rate and rhythm:

Atrial rate and rhythm:

PR interval:

QRS complex:

Identification:
Answer #1

Each large box in ECG is 0.2 second.Each small horizontal line is 0.04 second.

The R-R interval = 7 large boxes

So R-R interval = 7 * 0.2 = 1.4 second

So time difference between two ventricular beats (because QRS complex denotes ventricular contractions) = 1.4 second

Ventricular rate = 60/1.4 = around 43 beats/minute

Rhythm is regularly regular.

Atrial rate is also same because the time interval between 2 P waves is also 7 large boxes.

Rhythm is regularly regular.

PR interval = 4 small boxes = 0.04 * 4 = 0.16 seconds

QRS complex is around 2 small boxes = 0.04 *2 = 0.08 seconds

So this ECG indicates bradycardia (heart rate less than 60 beats per minute ).